### v1

# Enhancing Oscillator-Based Ising Machine Models with Amplitude Dynamics and Polynomial Interactions

Abstract
We present an oscillator model with both phase and amplitude dynamics for oscillator-based Ising machines that addresses combinatorial optimization problems with polynomial cost functions of arbitrary order. Our approach addresses fundamental limitations of previous oscillator-based Ising machines through a mathematically rigorous formulation with a well-defined energy function and corresponding dynamics. The model demonstrates monotonic energy decrease and reliable convergence to low-energy states. Empirical evaluations on 3-SAT problems show significant performance improvements over existing phase-amplitude models. Additionally, we propose a flexible, generalizable framework for designing higher-order oscillator interactions, from which we derive a practical method for oscillator binarization without compromising performance. This work strengthens both the theoretical foundation and practical applicability of oscillator-based Ising machines for complex optimization problems.
preprint: APS/123-QED
I INTRODUCTION
The limitations of traditional von Neumann architectures in efficiently tackling combinatorial optimization problems have spurred the exploration of alternative computing paradigms. Among these, Ising machines, inspired by the Ising model from statistical mechanics, have emerged as a promising approach. The versatility of Ising machines extends across a diverse range of applications, including portfolio optimization [1][2], drug discovery [3], vehicle routing [4, 5, 6], fraud detection [7], clustering [8], power grid optimization [9], wireless communication [10, 11], image restoration [12] and machine learning [13][14][15][16]. By mapping the decision variables of an optimization problem onto interacting spins [17], Ising machines can efficiently search for solutions by allowing the system to naturally evolve towards its lowest energy state. This process corresponds to finding a configuration of spins that minimizes the system’s energy, which directly translates to a high-quality solution to the original optimization problem. This approach offers the potential to significantly accelerate the solution of complex optimization problems that are intractable for classical computers.
A major challenge in Ising machine implementations is that the original Ising model supports only up to second-order interactions. Mapping practical problems such as 3-SAT introduces third-order terms into the Hamiltonian, creating a fundamental mismatch between the problem structure and the physical implementation. A widely adopted approach to address this issue is cubic-to-quadratic reduction [18, 19, 20, 21], which introduces auxiliary variables. However, this method often complicates the energy landscape, making it more difficult to navigate [22], and consequently leads to suboptimal performance [23, 24]. Recent advances [23, 25, 24, 26] have proposed native supports for higher-order interactions, avoiding the need for any reductions.
Ising machines based on networks implementing the Kuramoto model [27] of synchronization have gained significant traction. These oscillator-based Ising machines (OIMs) [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38] based on Kuramoto model exploit the emergent collective behavior of synchronized oscillators to emulate spin interactions, with the system’s natural tendency towards stable, synchronized states corresponding to low-energy solutions of the Ising Hamiltonian. While Kuramoto-based OIMs have shown promise, oscillator models that incorporate amplitude dynamics in addition to phase [23, 39] have aimed to provide a more complete physical description of the underlying hardware and improved solution quality.
With both phase and amplitude dynamics, these models better reflect the physical behavior of real-world oscillators and offer greater flexibility in oscillator network design [40]. For clarity, we refer to models that integrate both amplitude and phase in the context of Ising machines as the Hopf oscillator model.
Our study is motivated by the need for a faithful, general oscillator model that incorporates both phase and amplitude dynamics while supporting higher-order interactions. We begin by identifying critical limitations in the previously proposed Hopf oscillator model for optimization problems [23]:
  • Lack of a Physically Meaningful Hamiltonian Definition: A central concept in Ising machines is the system’s energy, which should correspond to the objective function being optimized. In previous work, the defined energy function is generally complex-valued, lacking a clear physical interpretation. This contrasts with traditional Ising machines where energy is a real-valued function directly related to the optimization problem cost function.
  • Non-Monotonic Energy Decrease: Ideally, the system’s energy should decrease monotonically as it evolves towards a solution. However, we observed that even the real part, imaginary part, and absolute value of the previously defined complex-valued energy in the Hopf oscillator model do not consistently decrease during the optimization process. This behavior violates a fundamental principle of Ising machines and hinders both theoretical understanding and practical performance.
To address these limitations, we have developed an oscillator model with complex-valued oscillator representation, based on Andronov-Hopf dynamics [40], providing a rigorous framework for oscillator-based Ising machines. Our key contributions include:
  • Energy Redefinition: We introduce the complex conjugate of the oscillator representation variable $z$ to define a real-valued and physically meaningful energy function that directly corresponds to the objective function of the optimization problem.
  • Oscillator Dynamics Derivation: Using Wirtinger calculus [41], we derive the oscillator dynamics from our redefined energy function. These dynamics guarantee a monotonic decrease in energy under ideal conditions as the system evolves towards a solution.
  • Generalization to Higher-Order Interactions: Our formulation naturally extends to include higher-order interactions among oscillators, thus broadening its applicability to more complex optimization problems.
  • Alternative Binarization Approach: We propose an alternative method for binarizing the oscillator variable $z$, derived from the model formulation. This approach maintains solution quality while providing flexibility in implementation.
This rigorous framework not only provides a deeper theoretical understanding of OIMs but also leads to significant empirical performance improvements compared to existing phase-amplitude approaches. Remarkably, the energy function terms in our model closely resemble those found in nonlinear optical or four-wave mixing Hamiltonian [42, 43]. This connection suggests exciting new paths for realizing powerful Ising machines, potentially leveraging existing photonic technologies. Furthermore, we demonstrate that our model reduces to a previously proposed high-order, Kuramoto-based formulation [26] once amplitude dynamics are removed. Our work therefore generalizes previous polynomial frameworks, offering perspective on the relationship between Hopf and Kuramoto oscillator models for Ising machines.
Our findings offer a stronger theoretical foundation for OIMs while providing actionable guidelines for leveraging their potential to address complex optimization challenges.
II BACKGROUND AND PRELIMINARIES
II.1 Ising Hamiltonian for Combinatorial Optimization
1-D spin glass systems were originally proposed as a model for quenched disorder and frustration in metallic alloys [44, 45]. For $N$ 1-D spins $s=(s_{1},...,s_{N})$ with a coupling matrix $J\in\mathbb{R}^{N\times N}$ and external bias $h\in\mathbb{R}^{N}$, the Ising spin glass (ISG) Hamiltonian, named after Ernst Ising for his earlier work on purely ferromagnetic systems [46], is given by
$H(s)=-\sum_{i<j}J_{ij}s_{i}s_{j}-\sum_{i=1}^{N}h_{i}s_{i}.$ (1)
Later work revealed that finding the ground state of a spin glass is NP-hard [47]. Accordingly, a number of high-impact combinatorial problems can be mapped to an equivalent spin glass system [17, 48]. In particular, an ISG Hamiltonian can be trivially mapped to Quadratic Unconstrained Binary Optimization (QUBO) problems, a well-studied class within operations research [49].
The relative simplicity of the ISG Hamiltonian makes it highly amenable for hardware acceleration, leading to the advent of “Ising machines” (IMs) [48]: devices which “implement” the Ising Hamiltonian in some manner. IM proposals include purely digital accelerators [50, 51], analog devices evolving a continuous relaxation of the Ising Hamiltonian [28, 52, 53], and hybrid digital/analog systems [54, 55].
II.2 Solving 3-SAT problems with Ising Machine
Ising machines have been successfully applied to quadratic unconstrained binary optimization (QUBO) problems [52, 28, 56] and even polynomial unconstrained binary optimization (PUBO) [26, 57, 23] problems with the help of higher-order interactions. A generic Ising machine workflow usually involves ① converting the problem to a discrete Ising formula, ② mapping the formula to a continuous-time, dynamical Ising machine capable of seeking the minimum energy state and ③ extracting the solution from the final device state. In this subsection, we demonstrate how to map 3-SAT problems to an Ising Machine as a preliminary step.
Consider a Boolean formula $f:\{0,1\}^{N}\to\{0,1\}$ in 3-literal conjunctive normal form (3-CNF)
$f(x)=\bigwedge_{j=1}^{m}(\ell_{j,1}\lor\ell_{j,2}\lor\ell_{j,3}),$ (2)
where each $\ell_{j,1}$ is a literal (either a variable $x_{p}$ or its negation $\overline{x}_{p}$) and each disjunction of literals $(\ell_{j,1}\lor\ell_{j,2}\lor\ell_{j,3})$ is a clause. The 3-SAT problem consists of finding some $x$ such that $f(x)=1$, i.e., an assignment satisfying all of the clauses. We restrict ourselves to 3-SAT without loss of generality, as any $K$-SAT problem can be reduced to an equivalent 3-SAT problem by adding $O(K\cdot m)$ variables and clauses [58].
Let $s=\{s_{i}\}_{i=1}^{n}\in\{-1,+1\}^{n}$ be the vector of Ising spins. We follow the standard convention that $s_{i}=1\iff x_{i}=1$ and $s_{i}=-1\iff x_{i}=0$ [17].
For each literal $\ell_{j,k}\in\{x_{p},\overline{x}_{p}\}$ we define a spin function $L_{j,k}:\{-1,1\}^{n}\to\{0,1\}$ as
$L_{j,k}(s)=\frac{1}{2}\begin{cases}1-s_{p}&\text{ if }\ell_{j,k}=x_{p}\\ 1+s_{p}&\text{ if }\ell_{j,k}=\overline{x}_{p}\end{cases},$ (3)
where $L_{j,k}(s)=0$ if the literal is satisfied, and $1$ otherwise.
A clause $c_{j}=(\ell_{j,1}\lor\ell_{j,2}\lor\ell_{j,3})$ is satisfied if at least one of its literals is true. Similar to individual literals, we can define a spin function $C_{j}(s)$ as
$C_{j}(s)=L_{j,1}(s)\cdot L_{j,2}(s)\cdot L_{j,3}(s),$ (4)
where $C_{j}(s)=0$ if $c_{j}$ is satisfied and 1 otherwise.
Substituting Eq. (3), we can see that each $C_{j}$ is a spin polynomial of degree three
$C_{j}(s)=\left(\frac{1\pm s_{p}}{2}\right)\left(\frac{1\pm s_{q}}{2}\right)\left(\frac{1\pm s_{r}}{2}\right),$ (5)
where $s_{p}$, $s_{q}$, and $s_{r}$ are spins corresponding to the variables of each literal.
We then define the objective function $E:\{-1,1\}^{N}\to\mathbb{Z}_{+}$ as the sum of all clause functions $C_{j}(s)$,
$E(s)=\sum_{j}C_{j}(s)=\sum_{j}L_{j,1}(s)L_{j,2}(s)L_{j,3}(s).$ (6)
From Eq. (5), when a clause is satisfied, its corresponding $C_{j}$ evaluates to zero, contributing nothing to the total energy. In contrast, if a clause is unsatisfied, $C_{j}$ contributes a positive value of $1$ to $\mathcal{H}$. Therefore, minimizing $\mathcal{H}$ effectively minimizes the number of unsatisfied clauses. The optimization problem can thus be formally stated as
$E$ $=\sum_{i=1}^{n}h_{i}s_{i}+\sum_{i=1}^{n}\sum_{j=1}^{n}J_{ij}s_{i}s_{j}$ (7)
$+\sum_{i=1}^{n}\sum_{j=1}^{n}\sum_{k=1}^{n}P_{ijk}s_{i}s_{j}s_{k}+\text{constant},$
where $\mathbf{h}$, $\mathbf{J}$, and $\mathbf{P}$ are the coefficient vector, the coupling matrix, and the three-body interaction tensor, respectively. Minimizing $E$ corresponds to finding an assignment of spin variables $s_{i}$ that satisfies the maximum number of clauses. A solution with $E=0$ indicates that all clauses are satisfied, thereby providing a solution to the original 3-SAT problem.
In practice, analog Ising machines often represent binary spins $s$ by some physical quantities, thus relaxing the discrete phase space into a continuous one. We assume native device support for 3-body interactions and focus on the performance of the specific formulation. 3-body interactions can be implemented in analog electronic systems by means of multi-input gates (see Refs. [26, 24] for more details).
II.3 Kuramoto Oscillator Model
The Kuramoto model [27] was originally proposed as a phase reduction method to study synchronization phenomena in weakly-coupled oscillator networks [59, 60]. A system of $N$ coupled oscillators with phase $\theta=(\theta_{1},...,\theta_{N})$ is described by a system of coupled ODEs
$\frac{d\theta_{i}}{dt}=\kappa\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}J_{ij}\sin(\theta_{i}-\theta_{j})+\omega_{i},$ (8)
where $\kappa$ and $J_{ij}$ describe pairwise and global interaction strengths respectively, and $\omega_{i}$ is the natural frequency of the $i^{th}$ oscillator. We set $\omega_{i}=0$ in Eq. (8) by assuming all oscillators have identical $\omega_{i}$ and viewing the system in a rotating reference frame. Using spin glass interaction coefficients for $J$ has led to oscillator-based accelerators for combinatorial optimization: oscillator Ising machines (OIM) [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 61, 62].
One can trivially derive the global Lyapunov function associated with Eq. (8) as
$E(\theta)=\kappa\sum_{i<j}J_{ij}\cos(\theta_{i}-\theta_{j}).$ (9)
The coupled oscillator system tends to minimize energy $E(\theta)$ over time [63], laying the foundation for OIMs that tackle combinatorial optimization problems. The original OIM architecture [28] utilized analog electronic LC-tank oscillators, and more recent proposals have broadened the choice of substrate to ring oscillators [64, 65], phase transition nano-oscillators [66, 67, 68], and magnetic tunnel junctions [69, 70].
II.4 Sub-harmonic Injection Locking
[…]
$\theta_{1}-\theta_{2}=\pi.$ (10)
For a toy 2-oscillator system, the binarization is clear. However, coupled Kuramoto oscillator networks may not permit a closed-form binarization scheme. Kuramoto synchronization results in the oscillators bifurcating to aligned/non-aligned phases along some axis in the imaginary plane; however, that axis is not known a priori. To provide a simple, bijective mapping between oscillator fixed points and spin states, techniques such as sub-harmonic injection locking (SHIL) are introduced.
To provide simple binarization, previous work [28] injects a sub-harmonic injection locking (SHIL) term to their spin dynamics,
$\frac{d\theta_{i}}{dt}=\kappa\sum_{i<j}^{N}J_{ij}\sin(\theta_{i}-\theta_{j})+\kappa_{S}(t)\overbrace{\sin(2\theta_{i})}^{\text{SHIL}}.$ (11)
The second harmonic creates a potential well at $\theta_{i}=n\pi$ for $n\in\mathbb{N}$, inducing sub-harmonic injection locking (SHIL). During annealing, periodic application of SHIL forces the oscillators to bifurcate, with appropriate hyperparameters ensuring a final two-cluster bifurcation.
III METHODS
III.1 Hopf Oscillator Model
While the Kuramoto model is sufficient to study phase-only phenomena, real-world oscillators will have variability in both phase and amplitude. To enable a more faithful model, we can generalize from the phase-reduced representation to complex $z\in\mathbb{C}^{N}$. An oscillator model for combinatorial optimization and sampling is the Adronov-Hopf model (or “Hopf model” for brevity) [23, 40]. The evolution of each Hopf oscillator is given by a system of coupled ODEs
$\frac{dz_{i}}{dt}=f(z_{i})+g(z_{i}),$ (12)
where $z_{i}$ describes the phase and amplitude of the $i$-th oscillator, $f(z_{i})$ represents the local dynamics of the oscillator, and $g(z_{i})$ represents the coupling interaction between oscillators. In previous work [23] (which is also the baseline we compare with) , the local dynamics $f(z_{i})$ is defined as
$f(z_{i})=(\lambda_{i}+i\omega_{i})z_{i}+\rho_{i}z_{i}|z_{i}|^{2},$ (13)
where the center frequency is set to $\omega_{i}=0$ for all oscillators. The parameters $\lambda_{i}$ and $\rho_{i}$ are chosen such that the system exhibits a stable limit cycle with unit amplitude. The term $g(z_{i})$ is derived from the model energy function $\mathcal{H}$. Previous work [23] directly maps the optimization problem to the system energy function $\mathcal{H}$ by replacing the Ising spin variable $s_{i}$ in Eq. (7) with the complex oscillator representation $z_{i}$ as follows:
$\mathcal{H}$ $=\sum_{i=1}^{n}h_{i}z_{i}+\sum_{i=1}^{n}\sum_{j=1}^{n}J_{ij}z_{i}z_{j}$ (14)
$+\sum_{i=1}^{n}\sum_{j=1}^{n}\sum_{k=1}^{n}P_{ijk}z_{i}z_{j}z_{k}+\text{constant}.$
Although $\mathcal{H}$ can be interpreted as a function of oscillator states, several observations motivate us to investigate a revised formulation:
  1. 1.
    The energy $\mathcal{H}(z)$ is generally complex-valued. While the function is holomorphic and mathematically differentiable, the physical interpretation of a complex energy remains ambiguous.
  2. 2.
    The model does not ensure that the real part, imaginary part, or absolute value of $\mathcal{H}(z)$ decreases monotonically under ideal conditions. This raises concerns about convergence and optimization efficiency. Lyapunov stability analysis, typically used to demonstrate that a system seeks a local minimum of an energy function, is not applicable here.
To address the challenges in accurately describing the system and the oscillator dynamics with a complex variable $z_{i}$,
  1. 1.
    We introduce the complex conjugate $z_{i}^{*}$ into the energy function. This approach enables the formulation of a real-valued energy function $\mathcal{H}^{\prime}(z)$ that provides a meaningful physical interpretation of the system.
  2. 2.
    We derive the oscillator dynamics from the redefined energy and include the negative complex conjugate of the gradient of the energy function. This adjustment naturally drives the system toward the minima of the real-valued energy function.
Specifically, the energy function of our Hopf oscillator model is defined as follows:
$\mathcal{H^{\prime}}$ $=\frac{1}{2}\sum_{i=1}^{n}h_{i}(z_{i}+z_{i}^{*})+\sum_{i=1}^{n}\sum_{j=1}^{n}J_{ij}z_{i}z_{j}^{*}$ (15)
$+\frac{1}{2}(\sum_{i=1}^{n}\sum_{j=1}^{n}\sum_{k=1}^{n}P_{ijk}z_{i}z_{j}z_{k}^{*}$
$+\sum_{i=1}^{n}\sum_{j=1}^{n}\sum_{k=1}^{n}P_{ijk}z_{i}z_{j}^{*}z_{k}^{*})$
[…]
Note that this construction is not unique; in Sec. IV.2 we discuss a general framework for generating appropriate energy functions. With the introduction of the complex conjugate, the energy function is no longer holomorphic. However, by applying Wirtinger calculus [41], we extend the concept of differentiation to the complex domain, allowing us to rigorously derive the system dynamics. We can treat the complex function $f(z)$ as a function of both $z$ and its conjugate $z^{*}$. For non-holomorphic functions, this requires differentiation with respect to both variables. The symmetry of our energy function with respect to $z$ and $z^{*}$ permits a simplification: we can compute the gradient by considering only the derivative with respect to either $z$ or $z^{*}$. Now we have,
$\frac{d\mathcal{H^{\prime}}}{dz_{i}}$ $=\frac{1}{2}h_{i}+\sum_{j=1}^{n}J_{ij}z_{j}^{*}$ (16)
$+\frac{1}{2}(\sum_{j=1}^{n}\sum_{k=1}^{n}P_{ijk}z_{j}z_{k}^{*}+\sum_{j=1}^{n}\sum_{k=1}^{n}P_{ijk}z_{j}z_{k}^{*}$
$+\sum_{j=1}^{n}\sum_{k=1}^{n}P_{ijk}z_{j}^{*}z_{k}^{*})+\text{constant}.$
To drive the system towards minimizing the energy, we set
$\frac{dz_{i}}{dt}=-(\frac{d\mathcal{H^{\prime}}}{dz_{i}})^{*}$ (17)
and therefore obtain
[…]
with equality if and only if $\frac{d\mathcal{H^{\prime}}}{dz_{i}}=0$. That is, at a stationary point. Let $B_{r}(z^{\prime})$ be the ball of radius $r$ centered at $z^{\prime}$. If the stationary point $z^{\prime}$ is a local minimum, then the Hessian $\nabla^{2}\mathcal{H^{\prime}}(z^{\prime})$ is positive definite. From the definition of the Hessian in Wirtinger calculus [41] and the form of $\mathcal{H^{\prime}}$, we can conclude that $\nabla^{2}\mathcal{H^{\prime}}(z^{\prime})$ is continuous with respect to its arguments. Therefore, there exists an $\epsilon$ such that for all $y\in B_{\varepsilon}(z^{\prime})$, $\nabla^{2}\mathcal{H^{\prime}}(y)$ is also positive definite. Therefore, the dynamics (18) are locally Lyapunov stable within some neighborhood for all minima.
III.2 Compatibility with Kuramoto model
The Kuramoto model was originally derived as a phase reduction for networks of weakly coupled oscillators [27]. Both the phase-only model (including the quadratic Kuramoto model) and the Hopf model only approximate real-world oscillators. However, the Hopf model is more expressive, as it accounts for amplitude variations. In the case of amplitude homogeneity, however, we expect that the Hopf model reduces to the phase-only model. Here, we show that the Hamiltonian of our model, Eq. (15) can be used to derive phase-only models found in previous literature in the context of cubic forms. In so doing, we demonstrate our model ① conforms to physical intuition and ② generalizes previous work.
Expressing $z_{i}=|z_{i}|e^{-i\theta_{i}}$, we have the equivalent expression
$\mathcal{H^{\prime}}(z)$ $=\frac{1}{2}\sum_{i=1}^{n}h_{i}|z_{i}|(e^{i\theta_{i}}+e^{-i\theta_{i}})$ (19)
$+\sum_{i=1}^{n}\sum_{j=1}^{n}J_{ij}|z_{i}||z_{j}|e^{i(\theta_{i}-\theta_{j})}$
$+\frac{1}{2}(\sum_{i=1}^{n}\sum_{j=1}^{n}\sum_{k=1}^{n}|z_{i}||z_{j}||z_{k}|e^{i(\theta_{i}+\theta_{j}-\theta_{k})}$
$+\sum_{i=1}^{n}\sum_{j=1}^{n}\sum_{k=1}^{n}P_{ijk}|z_{i}||z_{j}||z_{k}|e^{-i(\theta_{i}+\theta_{j}-\theta_{k})})$
[…]
Since $\mathcal{H^{\prime}}(z)$ is real by construction, we have
$\mathcal{H^{\prime}}(z)$ $=\operatorname{Re}[\mathcal{H^{\prime}}(z)]$ (20)
$=\sum_{i=1}^{n}h_{i}|z_{i}|\cos(\theta_{i})$
$+\sum_{i=1}^{n}\sum_{i=j}^{n}J_{ij}|z_{i}||z_{j}|\cos(\theta_{i}-\theta_{j})$
$+\sum_{i=1}^{n}\sum_{j=1}^{n}\sum_{k=1}^{n}|z_{i}||z_{j}||z_{k}|\cos(\theta_{i}+\theta_{j}-\theta_{k})$
[…]
If we assume amplitude homogeneity ($|z_{i}|=|z_{j}|=|z_{k}|=1$), which will remove the amplitude dynamics from the model, then we obtain
$\mathcal{H^{\prime}}(z)=$ $\sum_{i=1}^{n}h_{i}\cos(\theta_{i})+\sum_{i=1}^{n}\sum_{i<j}^{n}J_{ij}\cos(\theta_{i}-\theta_{j})$ (21)
$+\sum_{i=1}^{n}\sum_{j=1}^{n}\sum_{k=1}^{n}\cos(\theta_{i}+\theta_{j}-\theta_{k})$
[…]
which is identical to the form proposed in Ref. [26].
IV Results
IV.1 Model Evaluation
We compare our proposed Hamiltonian (Eq. (15)) against the previously established baseline [23] (Eq. (14)), which also incorporates the amplitude dynamics for oscillator-based Ising machines. We simulate these models by numerically solving their governing differential equations.
Overall Performance We evaluate model performance for cubic problems on Boolean Satisfiability Problems (SAT) from the SATLIB benchmark [71], using instances from the uf-20, uf-50, uf-75, uf-100 and uf-150 categories, totaling 1,700 instances. For each instance, 100 random initial conditions are generated.
Figure 1: Percentage of solvable instances across problem sets (UF-20 to UF-150) from SATLIB [71] under 100 runs per instance. The proposed Hopf model consistently outperforms the baseline model [23].
The results show that our approach consistently outperforms the baseline Hopf model, solving a higher percentage of instances and achieving lower final energy values more frequently. These improvements indicate superior optimization performance, as illustrated in Fig. 1 and Fig. 2.
Figure 2: Final energy distribution for solving 3-SAT problems across approximately 170,000 runs per model. The proposed Hopf model achieves lower final energies with higher frequency, demonstrating better performance in finding optimal or near-optimal solutions.
Energy minimization tendency Our proposed Hopf model has a clear real-valued Lyapunov function indicating the energy of the system. The derived oscillator dynamics also provides a clear energy minimization tendency.
Consider a simple example with three oscillators and the optimization target is $E=s_{1}s_{2}s_{3}$ where $s_{i}\in\{-1,+1\}$. Map the optimization problem to both of the models. As shown in Fig. 3, we observe distinct behaviors between the models. Neither the real part, imaginary part, nor the absolute value of the baseline Hopf oscillator model $\mathcal{H}$ decreases monotonically, while our proposed model will consistently seek the minimum of the energy.
Figure 3: Energy trajectories over time for the baseline model [23] and the proposed Hopf model.
IV.2 Higher-order Models with Wirtinger Potentials
For effective optimization of combinatorial problems and accurate representation of the Ising formulation, the energy function of an oscillator system ideally satisfies two key conditions:
  1. 1.
    It should establish a clear correspondence between the continuous system energy and the original optimization problem. One approach to achieve this is through a real-valued energy function.
  2. 2.
    Upon binarization, the energy function would optimally yield the same values as the discrete Ising formulation. For instance, with a complex oscillator representation $z$, one may map $\operatorname{Re}(z)>0$ to $s_{i}=+1$ and $\operatorname{Re}(z)<0$ to $s_{i}=-1$; alternatively, in a phase-only model, mapping $\phi_{i}=0$ to $s_{i}=+1$ and $\phi_{i}=\pi$ to $s_{i}=-1$ achieves this equivalence.
There are multiple ways to construct an energy function that satisfies these conditions. For simplicity, we focus on an oscillator system that involves only $m$ interactions of order $k$, assuming that all coupling coefficients are unity. In this case, the energy function can be expressed as follows:
$\mathcal{H}^{C}=\sum_{i=1}^{m}\prod_{j=1}^{k}\left(z^{i}_{j}+\left(z^{i}_{j}\right)^{*}\right)$ (22)
where $z^{i}_{j}$ denotes the complex representation of the $j$-th oscillator in the $i$-th interaction. We refer to this as the Complete Potential, or C-potential for short. However, a suitable subset of its terms can be selected to form a new energy function that still satisfies the desired conditions. When expanded, the products can be organized into $k+1$ groups of the form
$\mathcal{H}^{p}=\sum_{i}\left[\prod_{j=1}^{p}z^{i}_{j}\right]\left[\prod_{j=p+1}^{k}(z^{i}_{j})^{*}\right]$ (23)
where $p\in\{0,1,\dots,\lfloor{k/2}\rfloor\}$. Each $H^{p}$ is real-valued and has identical values for $\vec{z}\in\mathbb{R}^{n}$, preserving Property 1. We refer to the family of cost functions $\{\mathcal{H}^{p}\}$ as Wirtinger potentials for $\mathcal{H}$, with $\mathcal{H}^{p}$ referred to as the $p$-potential for $\mathcal{H}$.
For example, when $k=3$ and assuming unity coupling coefficients, the C-potential is
$\mathcal{H}^{C}=\sum_{i,j,k}\left(z_{i}+z_{i}^{*}\right)\left(z_{j}+z_{j}^{*}\right)\left(z_{k}+z_{k}^{*}\right).$ (24)
The 0-potential is then
$\mathcal{H}^{0}=\sum_{i,j,k}\left(z_{i}z_{j}z_{k}+z_{i}^{*}z_{j}^{*}z_{k}^{*}\right),$ (25)
which is manifestly real and preserves the Ising equivalence upon binarization. Alternatively, one may define the 1-potential as
$\mathcal{H}^{1}=\sum_{i,j,k}\left(z_{i}z_{j}z_{k}^{*}+z_{i}^{*}z_{j}^{*}z_{k}\right),$ (26)
which similarly fulfills the required conditions.
Wirtinger potentials enable flexibility in modeling oscillator systems while maintaining consistency with the Ising paradigm. For example, in a 3-SAT problem, the optimization target comprises three types of terms: linear, quadratic, and cubic (corresponding to $k=1$, $2$, and $3$, respectively). This means that multiple oscillator modeling options exist for each type of term. The energy function Eq. (15) is one such option, and our empirical evaluations indicate that this combination delivers the best performance in our benchmarks. However, it may not be the optimal choice for all problem types. In the following section (IV.3), we discuss a real case that leverages the variety of these models to achieve effective binarization for oscillator-based Ising machines. A thorough analysis of higher-order systems would require extensive theoretical and experimental study, which is beyond the scope of this work. Future research may explore these formulations to further enhance the capabilities of oscillator-based Ising machines.
IV.3 Binarization Behavior: Quadratic Problems
Recall that reading out solutions from oscillator-based Ising machines (OIMs) requires converting each oscillator’s continuous phase (from 00 to $2\pi$) into discrete Ising spins ($+1$ or $-1$). Phase conversion typically involves introducing a binarization axis $\vec{b}$, which separates phases into two groups. Denoting $z=x+iy$ in vector notation as $\vec{z}=(x,y)^{T}$, we define the binarization function $f_{B}(z)=\text{sign}(\vec{z}\times\vec{b})$. The most convenient choice for $\vec{b}$ is typically the imaginary axis, hence we have $f_{B}(x+iy)=\text{sign}(x)$. Fig. 4 illustrates three typical binarization behaviors. A system that has not binarized with respect to $\vec{b}$ will give an incorrect solution $f_{B}(\vec{z})$. The system has still bifurcated, however the axis of bifurcation $\vec{\mathfrak{b}}$ is not parallel to $\vec{b}$, causing erroneous spin assignments. Systems which are binarized with respect to $\vec{b}$ can differ in the magnitude of binarization. A weakly binarized system is susceptible to misclassifications owing to readout error, while a strongly binarized system is more robust to classification error. In this section, we explore the implications of our Wirtinger potential formalism for phase binarization.
Whether $\vec{\mathfrak{b}}=\vec{b}$ depends on the problem formulation. Consider a quadratic Hamiltonian $\mathcal{H}(s_{1},s_{2})=s_{1}s_{2}$. $\mathcal{H}$ has three main Wirtinger potentials, the 0-potential $\mathcal{H}^{0}(z_{1},z_{2})$ , the 1-potential $\mathcal{H}^{1}(z_{1},z_{2})$ , and their convex combination, the “complete” C-potential $\mathcal{H}^{C}(z_{1},z_{2})$. Their explicit forms are given by
$\begin{split}\mathcal{H}^{0}(z_{1},z_{2})&=z_{1}z_{2}+z^{*}_{1}z^{*}_{2}\\ \mathcal{H}^{1}(z_{1},z_{2})&=z_{1}^{*}z_{2}+z_{1}z^{*}_{2}\\ \mathcal{H}^{C}(z_{1},z_{2})&=0.5(z_{1}z_{2}+z^{*}_{1}z^{*}_{2})\\ &+0.5(z_{1}^{*}z_{2}+z_{1}z^{*}_{2}).\end{split}$ (27)
Each potential gives identical values at $z_{1},z_{2}\in\{|z|e^{i0},|z|e^{i\pi}\}$ (typical binarization reference points). However, they do not behave identically under phase translations. Consider a global phase offset $\phi$. The 1-potential remains invariant, as
$\begin{split}\mathcal{H}^{1}(e^{i\phi}z_{1},e^{i\phi}z_{2})&=e^{i(\phi-\phi)}z_{1}^{*}z_{2}+e^{i(\phi-\phi)}z_{1}z^{*}_{2}\\ &=z_{1}^{*}z_{2}+z_{1}z^{*}_{2}=\mathcal{H}^{1}(z_{1},z_{2}).\end{split}$ (28)
The same is not true, however, of the 0-potential or C-potential, where
$\begin{split}\mathcal{H}^{0}(e^{i\phi}z_{1},e^{i\phi}z_{2})&=e^{i2\phi}z_{1}z_{2}+e^{-i2\phi}z_{1}^{*}z_{2}^{*}\\ &\neq\mathcal{H}^{0}(z_{1},z_{2})\end{split}$ (29)
and
$\begin{split}\mathcal{H}^{C}(e^{i\phi}z_{1},e^{i\phi}z_{2})&=\gamma e^{i2\phi}z_{1}z_{2}+0.5e^{-i2\phi}z_{1}^{*}z_{2}^{*}\\ &+0.5(z_{1}^{*}z_{2}+z_{1}z_{2}^{*})\\ &\neq\mathcal{H}^{C}(z_{1},z_{2})\end{split}$ (30)
in general. For instance, for $z_{1}=e^{i0}$, $z_{2}=e^{i\pi}$, $\phi=\pi/3$, we have
$\begin{split}\mathcal{H}^{0}(e^{i\phi}z_{1},e^{i\phi}z_{2})&=\frac{1}{2}[\cos(\pi+\pi/3)+\cos(\pi-\pi/3)]\\ &=-0.5\neq-1.0=\mathcal{H}^{0}(z_{1},z_{2}).\end{split}$ (31)
Investigating the Hessians provides more context. For convenience, we assume that the local confining potentials confine each oscillator to unit amplitude, allowing for a phase-reduced model
$\begin{split}\mathcal{H}^{0}(\theta_{1},\theta_{2})&=\cos(\theta_{1}+\theta_{2})\\ \mathcal{H}^{1}(\theta_{1},\theta_{2})&=\cos(\theta_{1}-\theta_{2})\\ \mathcal{H}^{C}(\theta_{1},\theta_{2})&=0.5\cos(\theta_{1}-\theta_{2})+(1-0.5)\cos(\theta_{1}+\theta_{2})\end{split}$ (32)
with Hessians
$\begin{split}\nabla^{2}\mathcal{H}^{0}(\theta_{1},\theta_{2})&=\begin{bmatrix}-\cos(\theta_{1}+\theta_{2})&-\cos(\theta_{1}+\theta_{2})\\ -\cos(\theta_{1}+\theta_{2})&-\cos(\theta_{1}+\theta_{2})\end{bmatrix}\\ \nabla^{2}\mathcal{H}^{1}(\theta_{1},\theta_{2})&=\begin{bmatrix}-\cos(\theta_{1}-\theta_{2})&\cos(\theta_{1}-\theta_{2})\\ \cos(\theta_{1}-\theta_{2})&-\cos(\theta_{1}-\theta_{2})\end{bmatrix}\\ \nabla^{2}\mathcal{H}^{C}(\theta_{1},\theta_{2})&=0.5\nabla^{2}\mathcal{H}^{0}(\theta_{1},\theta_{2})\\ &+0.5\nabla^{2}\mathcal{H}^{1}(\theta_{1},\theta_{2})\end{split}$ (33)
by the linearity of the Hessian operator.
$\nabla^{2}\mathcal{H}^{1}(\theta_{1},\theta_{2})$ and $\nabla^{2}\mathcal{H}^{0}(\theta_{1},\theta_{2})$ have eigenvalues $\{0,2\}$, however their null spaces $\mathcal{N}$ are disjoint. Specifically, $\mathcal{N}(\nabla^{2}\mathcal{H}^{0})=\text{Span}\{(1,-1)^{T}\}$, $\mathcal{N}(\nabla^{2}\mathcal{H}^{1})=\text{Span}\{(1,1)^{T}\}$. In other words, the 1-potential has zero curvature along the $(1,1)^{T}$ vector while being strongly convex along $(1,-1)$. In contrast, the 0-potential is strongly convex along $(1,1)^{T}$ while having zero curvature along $(1,-1)$. Neither potential is strongly convex in the neighborhood of their fixed points, causing each to have an uncountable number of solutions.
However, the convex combination $\nabla^{2}\mathcal{H}^{C}(\theta_{1},\theta_{2})$ is full-rank and positive definite in the neighborhood of all $\{0,\pi\}$ minima, meaning it guarantees descent to the unique phase minimizer within each neighborhood.
SHIL performs a similar function, adding the diagonal matrix
$\nabla^{2}\mathcal{H}^{\text{SHIL}}=\begin{bmatrix}-\sin(2\theta_{1})&0\\ 0&-\sin(2\theta_{2})\end{bmatrix}$ (34)
which makes $\nabla^{2}\{\mathcal{H}^{1}+\mathcal{H}^{\text{SHIL}}\}$ full rank and positive definite in the neighborhood of $\{0,\pi\}$. Therefore, we have two options for binarizing quadratic Hamiltonians:
  1. 1.
    Add $\mathcal{H}^{\text{SHIL}}$ to the problem Hamiltonian
  2. 2.
    Use $\mathcal{H}^{C}(z_{1},z_{2})$
Option 1 is typical in oscillator-based IMs [28, 66], however, our analysis offers an alternative. $\mathcal{H}^{C}(z_{1},z_{2})$ or, more generally, a convex combination of the 1-potential $\mathcal{H}^{1}(z_{1},z_{2})$ and the 0-potential $\mathcal{H}^{0}(z_{1},z_{2})$, creates a strongly convex function in the neighborhood of $\{0,\pi\}$ fixed points. In the next section, we explore this possibility to induce binarization in 3-SAT problems while maintaining high-quality solutions.
IV.4 Binarization Behavior: SAT Problems
Figure 4: Example phase plots for a non-binarized (left), poorly binarized (middle) and a well-binarized (right) system with axis of binarization $\vec{b}$.
Figure 5: Polar histograms of phase values for the different binarization strategies discussed.
Similar to the quadratic case, there are multiple choices to formulate a cubic oscillator model. Our empirical analysis on 3-SAT problems shows that the model defined by Eq. (15) attains the highest solution quality, but only weakly binarizes. Conversely, replacing the terms in Eq. (15) with the complete form in Eq. (22) yields a new energy function which strongly binarizes but suffers from inferior optimization performance. In our experiments, this variant solved 61.8%, 17%, 2%, and 0% of the instances for the uf-50, uf-75, uf-100, and uf-150 benchmarks, respectively, based on 100 independent trials per instance, which is similar to the baseline we compared with. This trade-off motivates the question: Can we design an oscillator network that simultaneously (i) achieves high solution quality, (ii) respects the Ising Hamiltonian, and (iii) yields binary spin values at convergence?
To address this, we examine the contributions of each term in the formulation Eq. (15) independently. Linear and cubic terms inherently attain weak binarization, as their Hessians have non-zero curvature along $(1)^{T},(1,1,1)^{T}$. The primary difference between the Hamiltonians lies with the quadratic terms. As we showed in the previous subsection, C-potential quadratic Hamiltonians are strongly convex in the vicinity of $\{0,\pi\}$ phase points, promoting strong binarization. In contrast, the 1-potential does not binarize, as the system is invariant to global phase shifts.
Figure 6: Cumulative distribution of final energy values for the different binarization strategies discussed.
Therefore, introducing a time-varying quadratic Wirtinger potential can provide strong binarization force while preserving the optimization performance of the oscillator network. More specifically, we can get a model with stronger binarization tendency by modifying the quadratic terms (order $k=2$) to be the complete form we previously defined in Eq. (22) as follows:
$\mathcal{H^{\prime\prime}}$ $=\frac{1}{2}\sum_{i=1}^{n}h_{i}(z_{i}+z_{i}^{*})$ (35)
$+\boxed{\frac{1}{4}\sum_{i=1}^{n}\sum_{j=1}^{n}J_{ij}(2z_{i}z_{j}^{*}+z_{i}z_{j}+z_{i}^{*}z_{j}^{*})}$
$+\frac{1}{2}\Bigl{(}\sum_{i=1}^{n}\sum_{j=1}^{n}\sum_{k=1}^{n}P_{ijk}z_{i}z_{j}z_{k}^{*}$
$\quad+\sum_{i=1}^{n}\sum_{j=1}^{n}\sum_{k=1}^{n}P_{ijk}z_{i}z_{j}^{*}z_{k}^{*}\Bigr{)}$
[…]
It is important to note that after binarization, both Hamiltonian (15) and Eq. (35), as well as any combination of them, yield identical values to the discrete Ising formulation, although binarization introduces a slight modification in the energy landscape. To demonstrate the effectiveness of our binarization strategies, we evaluated four scenarios on five uf-100 3-SAT instances:
  • None: The system evolves under Hamiltonian (15).
  • Static Potential: The system evolves under Hamiltonian (35).
  • Annealed SHIL: The system uses Hamiltonian (15) augmented with a SHIL signal, where the coefficient increases linearly.
  • Annealed Potential: The system starts with Hamiltonian (15) and linearly transitions to Hamiltonian (35) to achieve binarization.
Fig. 5 presents the distribution of the final oscillator phases. In the None scenario, the distributions are weakly peaked at 00 and $\pi$. However, a significant proportion of the phases are spread between the peaks, indicating little inherent binarization. In contrast, both the SHIL and Annealed strategies exhibit a strong binarization tendency.
The cumulative distribution plot of the final solution energy in Fig. 6 illustrates the overall performance (solution quality). The narrower figure of merit (FoM) should be just the probability of reaching the ground state (energy=0). However, we note that these are stochastic results and one way to better see through the underlying reality is to see the distribution of all solutions. (Another way is to show the narrow FoM over more test benchmarks.) In the cumulative distribution, the higher the curve the better. As we can see, the SHIL and Annealed strategies deliver performance that is similar to or slightly better than that achieved by the system evolving under Hamiltonian (15) alone. We hypothesize that the induced binarization acts as a form of stochastic perturbation, which may help the system escape local minima during the optimization process.
V DISCUSSION
Phase reduction approaches, including the Kuramoto model, fail to capture the full dynamics of oscillator-based Ising machines in real-world applications. We present a comprehensive Hopf oscillator model that explicitly incorporates amplitude dynamics, addressing this fundamental limitation. Our model demonstrates superior performance, consistently converging to lower energy states compared to phase-only models.
Our investigation focuses on model-level improvements, specifically examining the energy distribution when solving 3-SAT problems and quantifying the number of solvable problems without heuristic enhancements. While recent work [24, 57, 72] has shown that incorporating heuristic mechanisms can significantly boost solver performance, our fundamental model improvements provide a more robust foundation for such enhancements.
Incorporating amplitude dynamics introduces an additional degree of freedom. The use of a complex oscillator representation paves the way for further research. We interpret the local dynamics term $f(z_{i})$ in the oscillator model (Eq. (12)) not only as a mechanism to preserve physical oscillator properties but also as a means to improve the performance of the oscillator-based Ising machine. Here, we provide a perspective for future studies: oscillator-based Ising machines rely on subharmonic injection locking (SHIL) and nonlinearity to achieve better performance on MAX-CUT problems [28, 61], whereas Linear Ising Machines (LIMs) like BRIM [52] do not. By engineering the local dynamics, one can steer the oscillator-based Ising machine to mimic the dynamics of LIMs. For example, in quadratic MAX-CUT problems, the oscillator dynamics $f(z_{i})$ could be designed to preserve the real part of the oscillator representation $z_{i}$ while adjusting the imaginary part to maintain essential oscillator properties. This approach shifts the optimization focus from hard-to-implement system-level strategies to tuning inherent oscillator properties, although further experimental validation remains necessary.
VI CONCLUSION
Our work is driven by the need for a more accurate model for oscillator-based Ising machines. We observed several fundamental inconsistencies in existing phase-amplitude models. By addressing these issues, we propose a model that incorporates both phase and amplitude dynamics, features a physically meaningful Lyapunov function, and remains compatible with the widely used phase-only models in oscillator-based Ising machine implementations. Experiments show that our model outperforms existing phase-amplitude approaches in optimization performance and exhibits a correct energy minimization tendency. Furthermore, we propose a method to generalize the model for constructing higher-order Ising machines and an implementable approach for extracting solutions from the oscillators. These findings contribute to both the theoretical understanding and practical realization of oscillator-based Ising machines, offering valuable directions for future research and development in this promising field.